Integrate dx/(25x²-4)½ using trigonometric substitution
1
Expert's answer
2020-02-25T11:23:34-0500

Substitution:

"x=\frac{2}{5}\sec{\theta}"

Then "dx=\frac{2}{5}\frac{\sin{\theta}}{\cos^2{\theta}}d\theta"

"\int\frac{dx}{\sqrt{25x^2-4}}=\int\frac{2}{5}\frac{\sin{\theta}}{\cos^2{\theta}\sqrt{4\sec^2{\theta}-4}}d\theta=\frac{1}{5}\int\frac{\sin{\theta}}{\cos^2{\theta}\sqrt{\sec^2{\theta}-1}}d\theta="

"=\frac{1}{5}\int\frac{\sin{\theta}}{\cos^2{\theta}\tan{\theta}}d\theta=\frac{1}{5}\int\frac{d\theta}{\cos{\theta}}=\frac{1}{5}\int\sec{\theta}d\theta=\frac{1}{5}\ln|\sec{\theta}+\tan{\theta}|+C.\\
http://www.math.com/tables/integrals/more/sec.htm"

"\sec{\theta}=\frac{5}{2}x"

"x^2=\frac{4}{25}\frac{1}{\cos^2{\theta}}\\
x^2-\frac{4}{25}=\frac{4}{25}tan^2{\theta}\\
tan^2{\theta}=\frac{25}{4}x^2-1\\
tan{\theta}=\sqrt{\frac{25}{4}x^2-1}"

We have

"\int\frac{dx}{\sqrt{25x^2-4}}=\frac{1}{5}\ln|\frac{5}{2}x+\sqrt{\frac{25}{4}x^2-1}|+C."